Lito is 3 times Sandra's age. The difference of their ages is 51. Find their present ages.
I know that Sandra can be denoted as x and Lito as 3x but after this, I don't know how to proceed. If the problem is the sum of their ages is 51, then I know that it would be x + 3x = 51. But the problem is the difference of their ages is 51. What do I do in this type of question?

Answer by ankor@dixie-net.com(22740)

Lito is 3 times Sandra's age.
The difference of their ages is 51. Find their present ages.
:
I know that Sandra can be denoted as x and Lito as 3x but after this, I don't know how to proceed.
If the problem is the sum of their ages is 51, then I know that it would be x + 3x = 51. But the problem is the difference of their ages is 51. What do I do in this type of question?
:
You write the equation: 3x - x = 51
Subtract like terms
2x = 51
x = 51/2
x = 25.5 yrs is Sandra
then
3(25.5) = 76.5 is Lito's age
:
:
Usually these problems result in integers but this one doesn't. check it
76.5 - 25.5 = 51